OpenStudy (thesecret20111):

expand and simplify 4√6 (2√3-√6) Please show steps I want to learn how to do it myself :D

OpenStudy (espex):

\[4\sqrt{6}(2\sqrt{3} - \sqrt6) \rightarrow (4\sqrt{6} * 2\sqrt{3}) - (4\sqrt{6} * \sqrt{6})\] \[8\sqrt{6*3} - 4\sqrt{6*6} \rightarrow 8\sqrt{18} - 4\sqrt{36}\] \[8\sqrt{2*9} - 24 \rightarrow 24\sqrt{2} - 24 \rightarrow 9.94\]

OpenStudy (thesecret20111):

So are you finding a number that gives an exact square, then turning the numbers under the root signs to a multiple of that exact square?

OpenStudy (espex):

OR from \[8\sqrt{2*9} - 24 \rightarrow 8\sqrt{18} - 4\sqrt{36} \rightarrow \rightarrow 8\sqrt{2*9} - 24 \rightarrow 24\sqrt{2} - 24\rightarrow8(3\sqrt{2} - 3)\]

OpenStudy (espex):

Distributing the \[4\sqrt{6}\] then looking to pull the squares from under the radical.

OpenStudy (espex):

That should have read: \[8\sqrt{2*9} - 24\rightarrow 24\sqrt{2} - 24\rightarrow8(3\sqrt{2} - 3)\]

OpenStudy (thesecret20111):

oh ok I get it now :) Thanks so much!

OpenStudy (espex):

You're welcome.